# Tight Bounds for Schrödinger Potential Estimation in Unpaired Data Translation

### v2

@@ Abstract
Modern methods of generative modelling and unpaired data translation based on Schrödinger bridges and stochastic optimal control theory aim to transform an initial density to a target one in an optimal way. In the present paper, we assume that we only have access to i.i.d. samples from initial and final distributions. This makes our setup suitable for both generative modelling and unpaired data translation. Relying on the stochastic optimal control approach, we choose an Ornstein-Uhlenbeck process as the reference one and estimate the corresponding Schrödinger potential. Introducing a risk function as the Kullback-Leibler divergence between couplings, we derive tight bounds on generalization ability of an empirical risk minimizer in a class of Schrödinger potentials including Gaussian mixtures. Thanks to the mixing properties of the Ornstein-Uhlenbeck process, we almost achieve fast rates of convergence up to some logarithmic factors in favourable scenarios. We also illustrate performance of the suggested approach with numerical experiments.
@@ 1 Introduction
The Schrödinger bridge problem has emerged as a powerful framework in modern generative modelling and unpaired image-to-image translation. It formulates the task of transforming an initial distribution into a target one via a probabilistic coupling that is as close as possible—in the sense of relative entropy—to a prescribed reference dynamics. In the present paper, we assume that both initial and target distributions are absolutely continuous and denote their densities with respect to the Lebesgue measure by $\rho_{0}$ and $\rho_{T}$, respectively.
Let $\mathsf{q}_{t}(y\,|\,x),$ $t\in[0,T],$ denote the transition density of a time-homogeneous reference Markov process $X^{0}$ (typically a diffusion), and let $\pi^{0}$ be the joint distribution of $(X_{0}^{0},X_{T}^{0})$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ given by the initial distribution $\rho_{0}$ and the forward dynamics of the reference process. Then, the Schrödinger bridge problem seeks a coupling $\pi(x,y)$ between $\rho_{0}$ and $\rho_{T}$ that minimizes the Kullback-Leibler (KL) divergence
@@
Under mild regularity conditions, the minimizer $\pi^{*}$ of the variational problem (1) admits a specific form (see (Leonard, 2014, Theorem 2.4)):
$\pi^{*}(x,y)=\nu_{0}^{\star}(x)\,\mathsf{q}_{T}(y\,|\,x)\,\nu_{T}^{\star}(y),$ (2)
where $\nu_{0}^{\star}$ and $\nu_{T}^{\star}$ are positive functions referred to as Schrödinger potentials. These scaling functions modify the reference transition structure so that the resulting coupling satisfies the marginal constraints. From this coupling, one can construct the corresponding Schrödinger bridge process $X^{*}$, which is a time-inhomogeneous Markov process with dynamics governed by a drift-modified version of the reference process.
@@
which is studied quite well. However, there are two drawbacks. First, the entropic optimal transport formulation deals with scalar parameter $\sigma$ and does not take into account possible anisotropy of the data. Second, if the reference dynamics obeys (3), then the correlation between $X_{0}^{0}$ and $X_{T}^{0}$ decays at slow polynomial rate. In the context of image-to-image translation, this could mean that initial images have an overabundant and possibly negative influence on the final ones. As a consequence, a learner has to take larger values of $\sigma$ or $T$ to mitigate this effect. In this work, we take the Ornstein-Uhlenbeck (OU) process as the reference dynamics. To be more specific, we will consider the base process $X_{t}^{0}$ solving the SDE
@@
${\mathrm{d}}X_{t}^{*}=b^{*}(t,X_{t}^{*}){\mathrm{d}}t+\Sigma^{1/2}{\mathrm{d}}W_{t},\quad X_{0}^{*}\sim\rho_{0},$ (5)
@@
$b^{*}(t,x)=b(m-x)+\nabla\log\nu^{\star}_{t}(x),$ (6)
and $\nu^{\star}_{t}(x)$ denotes the time-evolved Schrödinger potential given by
$\nu^{\star}_{t}(x)=\int\nu^{\star}_{T}(y)\,\mathsf{q}_{T-t}(y\,|\,x)\,{\mathrm{d}}y.$
@@
$\varphi^{\circ}\in\operatornamewithlimits{argmin}\limits_{\varphi\in\Phi}{\mathsf{KL}}(\pi^{*},\pi^{\varphi}),$
where
${\mathsf{KL}}(\pi^{*},\pi^{\varphi})=\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\log\frac{\pi^{*}(x,y)}{\pi^{\varphi}(x,y)}\,\pi^{\varphi}(x,y)\,{\mathrm{d}}x\,{\mathrm{d}}y.$
@@
${\mathcal{T}}_{t}[g](x)=\int\limits_{\mathbb{R}^{d}}g(y)\,\mathsf{q}_{t}(y\,|\,x)\,{\mathrm{d}}y$
@@
$\pi^{*}(x,y)=\frac{\rho_{0}(x)\,\mathsf{q}_{T}(y\,|\,x)\,e^{\varphi^{*}(y)}}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](x)},\quad\text{where}\quad\varphi^{*}=\log(\nu^{*}_{T}).$ (8)
@@
The main theoretical contribution of this paper is a non-asymptotic high-probability generalization error bound for the empirical risk minimizer (11). Using concentration arguments and exploiting the ergodic properties of the OU process, we establish nearly fast convergence rates (up to logarithmic factors) for the population KL risk. These results highlight the statistical efficiency of the Schrödinger bridge approach in the finite-sample regime. Finally, we validate our theoretical findings with numerical experiments on synthetic and real data, demonstrating the robustness and flexibility of our method in practical generative modelling tasks.
Related work.
In (Rigollet and Stromme, 2025), the authors considered the following empirical dual objective:
$\widehat{S}(\nu_{0},\nu_{T}):=\frac{1}{n}\sum_{j=1}^{n}\log\nu_{0}(Z_{j})+\frac{1}{n}\sum_{i=1}^{n}\log\nu_{T}(Y_{i})-\frac{\sigma^{2}}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\nu_{0}(Z_{j})\,\mathsf{q}_{T}(Y_{i}\,|\,Z_{j})\,\nu_{T}(Y_{i})+\sigma^{2}.$ (13)
This objective arises as a variational representation of the entropic optimal transport problem, where the regularization term penalizes deviations from the empirical product measure $\rho_{0}^{n}\otimes\rho_{T}^{m}$, with regularization parameter $\sigma^{2}>0$. Note that since $\rho_{0}$ is fixed in our setting, the empirical KL objective (12) reduces to
@@
Let us also mention the work Pooladian and Niles-Weed (2024), where the authors show that the Schrödinger potentials obtained from solving the static entropic optimal transport problem between source and target samples can be modified to yield a natural plug-in estimator of the time-dependent drift defining the Schrödinger bridge between the two distributions. This construction related to the so-called Föllmer bridge provides a principled and computationally efficient way to approximate the dynamics of the bridge process directly from sample data, and complements our focus on coupling-level guarantees. However, the statistical guarantees established in their work are given in terms of the total variation distance between the laws of the estimated and the true Schrödinger processes, and are only valid on time intervals bounded away from the terminal time. The authors show that the error can grow as time approaches the final time point and that this growth rate depends critically on the intrinsic dimension of the problem. This highlights a fundamental challenge in learning Schrödinger bridges near the boundary of the time interval. The connection between the Schrödinger bridge (SB) problem and entropic optimal transport was also exploited in Tang et al. (2024), where the authors use a static entropy-regularized optimal transport formulation—along with a minibatch approximation-to efficiently learn the Schrödinger bridge without requiring simulation of the underlying stochastic process. This approach leverages the tractability of static OT to bypass the computational challenges associated with solving stochastic control problems, while still recovering meaningful couplings for generative tasks.
Let us stress that all the above approaches restrict their attention to the case where the reference process is Brownian motion. While this choice offers analytical convenience and aligns with the classical formulation of the Schrödinger problem, it limits modeling flexibility—especially in contexts where prior structure or time-inhomogeneity is important. In contrast, our framework accommodates more general reference processes, such as the Ornstein–Uhlenbeck process, enabling better control of prior dynamics and improved statistical regularity through mixing properties.
Notations.
The notation $f\lesssim g$ or $g\gtrsim f$ means that $f={\mathcal{O}}(g)$. Besides, we often replace $\max\{a,b\}$ and $\min\{a,b\}$ by shorter expressions $a\vee b$ and $a\wedge b$, respectively. For any $s\geqslant 1$, the Orlicz $\psi_{s}$-norm of a random variable $\xi\sim\mathsf{p}$ is defined as
$\|\xi\|_{\psi_{s}}=\inf\left\{u>0:\mathbb{E}e^{|\xi|^{s}/u^{s}}\leqslant 2\right\}.$
Sometimes, we use the notation $\|\xi\|_{\psi_{s}(\mathsf{p})}$ to emphasize that $\xi$ is drawn according to $\mathsf{p}$. Finally, given $p\geqslant 1$ and a probability density $\rho$, the weighted $L_{p}$-norm of a function $f$ is defined as $\|f\|_{L_{p}(\rho)}=(\mathbb{E}_{\xi\sim\rho}|f(\xi)|^{p})^{1/p}$. Given two probability densities $\mathsf{p}\ll\mathsf{q}$, the Kullback-Leibler divergence between them is defined as ${\mathsf{KL}}(\mathsf{p},\mathsf{q})=\mathbb{E}_{\xi\sim\mathsf{p}}\log\big(\mathsf{p}(\xi)/\mathsf{q}(\xi)\big)$.
@@ 2 Main Result
We do not require $\rho_{0}$ to be bounded away from zero or infinity on $\mathrm{supp}(\rho_{0})$. For this reason, Assumption 2 is satisfied for a large class of densities. A requirement on the target distribution is even milder, we just assume that it is sub-Gaussian.
@@
We would like to note that (14) yields that $\mathbb{E}_{Y\sim\rho_{T}}Y=0$. This is just a normalization condition, which does not diminish generality of the setup we consider. We proceed with two assumptions on properties of the log-potential $\varphi^{*}$ associated with the optimal coupling $\pi^{*}$ (see (8)) and the reference class $\Phi$.

### v3

@@ Abstract
Modern methods of generative modelling and unpaired data translation based on Schrödinger bridges and stochastic optimal control theory aim to transform an initial density to a target one in an optimal way. In the present paper, we assume that we only have access to i.i.d. samples from the initial and final distributions. This makes our setup suitable for both generative modelling and unpaired data translation. Relying on the stochastic optimal control approach, we choose an Ornstein-Uhlenbeck process as the reference one and estimate the corresponding Schrödinger potential. Introducing a risk function as the Kullback-Leibler divergence between couplings, we derive tight bounds on the generalization ability of an empirical risk minimizer over a class of Schrödinger potentials, including Gaussian mixtures. Thanks to the mixing properties of the Ornstein-Uhlenbeck process, we almost achieve fast rates of convergence, up to some logarithmic factors, in favourable scenarios. We also illustrate the performance of the suggested approach with numerical experiments.
@@ 1 Introduction
The Schrödinger bridge problem has emerged as a powerful framework in modern generative modelling and unpaired data translation (in particular, image-to-image translation). It formulates the task of transforming an initial distribution into a target one via a probabilistic coupling that is as close as possible to a prescribed reference dynamics in the sense of relative entropy. In the present paper, we assume that both the initial and target distributions are absolutely continuous and denote their densities with respect to the Lebesgue measure by $\rho_{0}$ and $\rho_{T}$, respectively.
Let $\mathsf{q}_{t}(y\,|\,x),$ $t\in[0,T],$ denote the transition density of a time-homogeneous reference Markov process $X^{0}$ (typically a diffusion), and let $\pi^{0}$ be the joint distribution of $(X_{0}^{0},X_{T}^{0})$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ given by the initial distribution $\rho_{0}$ and the forward dynamics of the reference process. Then, in the Schrödinger bridge problem one looks for a coupling $\pi(x,y)$ between $\rho_{0}$ and $\rho_{T}$ that minimizes the Kullback-Leibler (KL) divergence
@@
Under mild regularity conditions, the minimizer $\pi^{*}$ of the variational problem (1) admits a specific form (see Leonard (2014, Theorem 2.4)):
$\pi^{*}(x,y)=\nu_{0}^{*}(x)\,\mathsf{q}_{T}(y\,|\,x)\,\nu_{T}^{*}(y),$ (2)
where $\nu_{0}^{*}$ and $\nu_{T}^{*}$ are positive functions referred to as Schrödinger potentials. These scaling functions modify the reference transition structure so that the resulting coupling satisfies the marginal constraints. From this coupling, one can construct the corresponding Schrödinger bridge process $X^{*}$, which is a time-inhomogeneous Markov process with dynamics governed by a drift-modified version of the reference process.
@@
which is studied quite well. However, there are two drawbacks. First, the entropic optimal transport formulation deals with the scalar parameter $\sigma$ and does not take into account possible anisotropy of the data. Second, if the reference dynamics obeys (3), then the correlation between $X_{0}^{0}$ and $X_{T}^{0}$ decays at slow polynomial rate. In the context of image-to-image translation, this could mean that initial images have an overabundant and possibly negative influence on the final ones. As a consequence, a learner has to take larger values of $\sigma$ or $T$ to mitigate this effect. In this work, we take the Ornstein-Uhlenbeck (OU) process as the reference dynamics. To be more specific, we will consider the base process $X_{t}^{0}$ solving the SDE
@@
${\mathrm{d}}X_{t}^{*}=\beta^{*}(t,X_{t}^{*}){\mathrm{d}}t+\Sigma^{1/2}{\mathrm{d}}W_{t},\quad X_{0}^{*}\sim\rho_{0},$ (5)
@@
$\beta^{*}(t,x)=b(m-x)+\Sigma\nabla\log\nu^{*}_{t}(x),$ (6)
and $\nu^{*}_{t}(x)$ denotes the time-evolved Schrödinger potential given by
$\nu^{*}_{t}(x)=\int\nu^{*}_{T}(y)\,\mathsf{q}_{T-t}(y\,|\,x)\,{\mathrm{d}}y.$
@@
$\varphi^{\circ}\in\operatornamewithlimits{argmin}\limits_{\varphi\in\Phi}{\mathsf{KL}}(\pi^{*},\pi^{\varphi}),\quad\text{where}\quad{\mathsf{KL}}(\pi^{*},\pi^{\varphi})=\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\log\frac{\pi^{*}(x,y)}{\pi^{\varphi}(x,y)}\,\pi^{\varphi}(x,y)\,{\mathrm{d}}x\,{\mathrm{d}}y.$
@@
${\mathcal{T}}_{t}[g](x)=\int\limits_{\mathbb{R}^{d}}g(y)\,\mathsf{q}_{t}(y\,|\,x)\,{\mathrm{d}}y$ (7)
@@
$\pi^{*}(x,y)=\frac{\rho_{0}(x)\,\mathsf{q}_{T}(y\,|\,x)\,e^{\varphi^{*}(y)}}{{\mathcal{T}}_{T}[e^{\varphi^{*}}](x)},\quad\varphi^{*}=\log(\nu^{*}_{T}).$ (9)
@@
Table 1: Frequently used notations
Notation Eq. Meaning
$\rho_{0}$ Source (initial) distribution
$\rho_{T}$ Target (terminal) distribution
$\mathsf{q}_{t}(y\,|\,x)$ Transition density of the OU process, density of ${\mathcal{N}}(m_{t}(x),\Sigma_{t})$
$m_{t}(x)$ $(1-e^{-bt})m+e^{-bt}x$
$\Sigma_{t}$ $(1-e^{-2bt})\Sigma/(2b)$
$\nu_{0}^{*}$, $\nu_{T}^{*}$ (2) Left and right Schrödinger potentials
$\varphi^{*}$ (9) Target Schrödinger log-potential
$\pi^{*}$ (2), (9) Target coupling, solution of the variational problem (1)
$\pi^{\varphi}$ (8) Coupling corresponding to the log-potential candidate $\varphi$
$\widehat{\varphi}$ (12) Log-potential estimate, empirical risk minimizer
$\widehat{\pi}$ (12) Coupling estimate, corresponding to $\widehat{\varphi}$
@@
  • The main contribution of this paper is a non-asymptotic high-probability generalization error bound for the empirical risk minimizer (12) (Theorem 1). Its proof relies on concentration arguments and ergodic properties of the OU process.
  • When the approximation error is small, we establish nearly fast convergence rates (up to logarithmic factors) for the population KL risk. These results highlight the statistical efficiency of the Schrödinger bridge approach in the finite-sample regime.
  • We conduct numerical experiments on synthetic and real data illustrating that a proper choice of the reference process can improve data generation quality.
Notations.
The notation $f\lesssim g$ or $g\gtrsim f$ means that $f={\mathcal{O}}(g)$. Besides, we often replace $\max\{a,b\}$ and $\min\{a,b\}$ by shorter expressions $a\vee b$ and $a\wedge b$, respectively. For any $s\geqslant 1$, the Orlicz $\psi_{s}$-norm of a random variable $\xi\sim\mathsf{p}$ is defined as $\|\xi\|_{\psi_{s}}=\inf\left\{u>0:\mathbb{E}e^{|\xi|^{s}/u^{s}}\leqslant 2\right\}$. Sometimes, we use the notation $\|\xi\|_{\psi_{s}(\mathsf{p})}$ to emphasize that $\xi$ is drawn according to $\mathsf{p}$. Finally, given $p\geqslant 1$ and a probability density $\rho$, the weighted $L_{p}$-norm of a function $f$ is defined as $\|f\|_{L_{p}(\rho)}=(\mathbb{E}_{\xi\sim\rho}|f(\xi)|^{p})^{1/p}$. Given two probability densities $\mathsf{p}\ll\mathsf{q}$, the Kullback-Leibler divergence between them is defined as ${\mathsf{KL}}(\mathsf{p},\mathsf{q})=\mathbb{E}_{\xi\sim\mathsf{p}}\log\big(\mathsf{p}(\xi)/\mathsf{q}(\xi)\big)$. Some other frequently used notations are collected in Table 1.
2 Related work
The empirical success of denoising diffusion models (Ho et al., 2020; Song et al., 2021) and flow matching (Lipman et al., 2023) made researchers study more general frameworks for mapping an initial distribution to a target one, such as stochastic interpolants (Albergo & Vanden-Eijnden, 2023; Albergo et al., 2025), optimal transport (Tong et al., 2020; 2024b; 2024a) and Schrödinger bridges (De Bortoli et al., 2021; Shi et al., 2023; Liu et al., 2023). In the context of the Schrödinger bridge problem, a lot of attention was paid to development of solvers based on iterative proportional fitting (De Bortoli et al., 2021) and iterative Markovian fitting (Shi et al., 2023; Liu et al., 2023; Gushchin et al., 2024b). Some recent papers (Rapakoulias et al., 2025; Domingo-Enrich et al., 2025) suggest using stochastic optimal control theory (see Dai Pra (1991) for background) to construct Schrödinger bridges. Besides, a lot of efforts were made to construct efficient algorithms for solving the Schrödinger bridge problem and accelerate statistical inference (see, for instance, Tang et al. (2024); Gushchin et al. (2024a; b)). Despite numerous methodological advances in application of Schrödinger bridge framework to generative modelling, its theoretical justification is often quite challenging and largely missing. For instance, only recently Conforti et al. (2024) (see also Eckstein (2025)) extended results on geometric convergence of IPF from compactly supported $\rho_{0}$ and $\rho_{T}$ to strongly log-concave marginals with unbounded supports. We would like to emphasize that Conforti et al. (2024) consider an idealized setting ignoring sampling error. A comprehensive analysis of the statistical guarantees for the Schrödinger potential estimation is still to be done. We provide a detailed discussion for few contributions in this direction below.
In Rigollet & Stromme (2025), the authors considered the following empirical dual objective:
$\widehat{S}(\nu_{0},\nu_{T}):=\frac{1}{n}\sum_{j=1}^{n}\log\nu_{0}(Z_{j})+\frac{1}{n}\sum_{i=1}^{n}\log\nu_{T}(Y_{i})-\frac{\sigma^{2}}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\nu_{0}(Z_{j})\,\mathsf{q}_{T}(Y_{i}\,|\,Z_{j})\,\nu_{T}(Y_{i})+\sigma^{2}.$
This objective arises as a variational representation of the entropic optimal transport problem, where the regularization term penalizes deviations from the empirical product measure $\rho_{0}^{\otimes n}\otimes\rho_{T}^{\otimes n}$, with regularization parameter $\sigma^{2}>0$. Note that since $\rho_{0}$ is fixed in our setting, the empirical KL objective (13) reduces to
@@ 2 Related work
[…]
Let us also mention the work (Pooladian & Niles-Weed, 2025), where the authors show that the Schrödinger potentials obtained from solving the static entropic optimal transport problem between source and target samples can be modified to yield a natural plug-in estimator of the time-dependent drift defining the Schrödinger bridge between the two distributions. The authors show that, if both $\rho_{0}$ and $\rho_{T}$ are supported on compact $k$-dimensional submanifolds in $\mathbb{R}^{d}$, then it holds that
$\mathbb{E}\mathrm{TV}^{2}(\mathsf{P}_{[0,\tau]}^{*},\widehat{\mathsf{P}}_{[0,\tau]})\lesssim\frac{1}{\sqrt{n}}+\frac{1}{(T-\tau)^{k+2}n},\quad 0<\tau<T.$ (14)
Here $\mathsf{P}_{[0,\tau]}^{*}$ and $\widehat{\mathsf{P}}_{[0,\tau]}$ are the path measures corresponding to the optimally controlled process (5) and the process with estimated drift. Unfortunately, the right-hand side of (14) blows up as $\tau$ approaches TT [^1] Probably, this is due to singularity of the target distribution with respect to the Lebesgue measure in $\mathbb{R}^{d}$.. This highlights a fundamental challenge in learning Schrödinger bridges near the boundary of the time interval.
Let us stress that Korotin et al. (2024), Pooladian & Niles-Weed (2025), and Rigollet & Stromme (2025) restrict their attention to the case where the reference process is Brownian motion. While this choice offers analytical convenience and aligns with the classical formulation of the Schrödinger problem, it limits modelling flexibility, especially in contexts where prior structure or time-inhomogeneity is important. In contrast, our framework accommodates more general reference processes, such as the Ornstein–Uhlenbeck process, enabling better control of prior dynamics and improved statistical regularity through mixing properties.
@@ 3 Main Result
We do not require $\rho_{0}$ to be bounded away from zero or infinity on $\mathrm{supp}(\rho_{0})$. For this reason, Assumption 2 is satisfied for a large class of densities. It plays a crucial role in the proof of Lemma C.1, which helps us to verify a Bernstein-type condition (B.2). Extension of this result to the case of unbounded $\mathrm{supp}(\rho_{0})$ will require significant efforts. A requirement on the target distribution is even milder, we just assume that it is sub-Gaussian.
@@
We would like to note that (15) yields that $\mathbb{E}_{Y\sim\rho_{T}}Y=0$. This is just a normalization condition, which does not diminish generality of the setup we consider. In Korotin et al. (2024), the authors assumed that both $\rho_{0}$ and $\rho_{T}$ have compact supports. In contrast, we allow $\mathrm{supp}(\rho_{T})$ to be unbounded. We proceed with two assumptions on properties of the log-potential $\varphi^{*}$ associated with the optimal coupling $\pi^{*}$ (see (9)) and the reference class $\Phi$.